A company’s cost of equity is 21 percent and the before tax cost of debt is 14 percent. It has a

net worth of INR3400 crore and borrowings of INR1360 crore. The market capitalization of the

company is INR 5100 crore. The tax rate is 30 percent. What is the Weighted Average Cost of

Capital? What is a risk-free rate and its application in Valuing a Firm?



1
Expert's answer
2021-12-20T10:19:04-0500

WACC Estimation:

Cost of equity = 21 percent

Before tax cost of debt =14 percent

Tax rate = 30 percent

Amount of borrowings = INR1360 crore

Amount of equity = INR3400 crore


"WACC=\frac{E}{D+E}(r_e)+\frac{D}{D+E}(r_d)(1-t)"


"WACC=\frac{3400}{1360+3400}(0.21)+\frac{1360}{1360+3400}(0.14)(1-0.30)"


WACC = 12.8%


Risk-free rate and its application:

Risk-free rate means the minimum rate of return that an investor expects from any investment opportunity. The rate of return of an investment with zero risks is also referred as the risk-free rate. The yield on government bonds or treasury bills generally refers as the risk-free rate. It is because it is fixed rate of return for a set period of time.

The risk-free rate has greater application in valuing a firm. It is because it is useful to calculate the required rate of return under CAPM (capital asset pricing model), which is significant in the valuation by discounting free cash flows.